A mechanical toy launches a toy missile from the ground into the air at a velocity of 13 m/s at an angle of 58°. Determine the maximum height of the missile and how far the missile will land if the missile lands at the same level from which it was launched.


1
Expert's answer
2021-09-23T15:49:29-0400
"H=\frac{13^2\sin^2{58}}{2(9.8)}=6.2\ m\\R=\frac{13^2\sin{116}}{(9.8)}=15.5\ m"